# Stability of low-momentum rest frame modes in first- and second-order spin hydrodynamics

### v1

@@ III.6 Stability Conditions
To guarantee that perturbations of the equilibrium state do not grow exponentially with time, we demand that they satisfy the condition $\mbox{Im}[\omega(k_{z})]<0$. In the sound channels, this leads to the following list of conditions:
@@
IV Conclusion
@@
In this work we have analyzed the low-momentum rest frame stability of the second-order spin hydrodynamics. We have used the truncation scheme of the Israel-Stewart formalism derived in our earlier work that extends the minimal causal formulation studied earlier. The crucial property to achieve stability is the use of the spin equation of state, which allows for opposite signs of the electric and magnetic components of the spin density tensor. Our analysis holds for both first-order (Navier-Stokes-like) and optimal second-order dissipative (Israel-Stewart-like) spin hydrodynamics, i.e., the 40=16+24 relaxation-type dynamical equations. The results show that both frameworks are stable in the low-momentum regime, which is also in agreement with conventional relativistic hydrodynamics in this case. These findings establish a solid foundation for exploring causality and stability of the hydrodynamic frameworks with spin for different momentum ranges.
@@
Acknowledgements: A.D. thanks Eduardo Grossi for discussions related to the early stages of this study during the visit to INFN-Florence. A.D. gratefully acknowledges technical discussions with Tomoki Goda, Shi Pu, and Arpan Das. F.T. acknowledges the hospitality and support of the Institute of Theoretical Physics of the Jagiellonian University. This work was supported in part by the Polish National Agency for Academic Exchange NAWA under the Programme STER–Internationalisation of Doctoral schools, Project no. PPI/STE/2020/1/00020 (A.D.), and the Polish National Science Centre Grants No. 2018/30/E/ST2/00432 (R.R., A.D.) and 2022/47/B/ST2/01372 (W.F.).

### v2

@@ IV Stability Study
To guarantee that perturbations of the equilibrium state do not grow exponentially with time, we demand that they satisfy the condition
$\mbox{Im}[\omega(k_{z})]<0.$ (120)
IV.1 Low-momentum regime
In the sound channels, this leads to the following list of conditions:
@@ IV.1 Low-momentum regime
IV.2 High-momentum regime
To determine the range of applicability of our relativistic spin hydrodynamic model, it is crucial to check the behavior of the perturbations also in the $k_{z}\rightarrow\infty$ limit. To get stable perturbations in high-momentum limit, the modes (108)-(119) demand according to (120)
${\tau_{\Phi}>0,\quad\tau_{q}>0,\quad\tau_{\tau_{a}}>0,\quad\tau_{\tau_{s}}>0,\quad\tau_{\pi}>0,\quad\tau_{\Pi}>0,\quad\tau_{\phi}>0,\quad\gamma>0,\quad\tau_{\Theta}>0,\quad\tau_{q}>\lambda^{{}^{\prime}},}$ (132)
${\tilde{\chi}_{1}<0,\quad\tilde{\chi}_{2}<0,\quad\tilde{\chi}_{3}<0,}$ (133)
${c_{0}>0,\quad d_{0}>0,}$ (134)
${\frac{1}{\tau_{\pi}}+\frac{1}{\tau_{\phi}}+\frac{\lambda^{\prime}}{\tau_{q}(\tau_{q}-\lambda^{\prime})}-\frac{\gamma^{\prime}+\eta}{\gamma^{\prime}\tau_{\pi}+\eta\tau_{\phi}}>0}.$ (135)
Our verification process has demonstrably confirmed that all the above conditions are satisfied and exhibit complete congruence with the established conditions in the low-momentum limit, as detailed in IV.1.
This provides a compelling foundation for a conclusion: second-order spin hydrodynamics demonstrably exhibits stability. While our present investigation has prioritized the stability question within the high-momentum regime, there remains a compelling need for future studies to delve into the realm of stability analysis at finite momentum values. This deeper examination could be effectively accomplished by employing the well-established Routh-Hurwitz criterion [71, 86, 87]. By incorporating a more comprehensive analysis at finite momenta, we can ensure a more robust understanding of the system’s behavior.
V Asymptotic Causality Study
The causality of matter generally refers to its property of admitting only perturbations that propagate no faster than the speed of light in vacuum within the fluid (i.e., not admitting superluminal perturbations). After checking stability in section (IV), verifying causality is crucial. Without causality, any solutions to the equations may not have physical meaning. If our system satisfies causality, we can solve our set of 40 interconnected relaxation-type dynamical equations and describe physical observables. This would allow us in the future to connect our formalism to other relevant topics [88, 89, 90, 91]. Here, we follow the seminal work of [92], followed by [76], where the asymptotic causality conditions boil down to,
[…]
and
[…]
Note that, as formulated mathematically in [92] and illustrated physically in [76], the first condition alone is not sufficient to guarantee causality  [^3] Condition (136) is also written in some literature as $\lim_{k_{z}\to\infty}\frac{\partial Re\omega(k_{z})}{\partial k_{z}}\leq 1$. This can be verified as all the real parts of the dispersion relations in this limit approach a constant multiple of $k_{z}$.. This can be seen through the non-relativistic, acausal diffusion mode with $\omega\sim ik^{2}$, which satisfies equation (136) but not (137).
Therefore, we need to the following constraints:
${0\leq-\frac{\frac{4}{3}\tilde{\chi}_{2}\tau_{\Phi}+\tilde{\chi}_{1}\tau_{\tau_{s}}}{\tau_{\tau_{s}}\tau_{\Phi}}\leq 1,}$ (138)
${0\leq\frac{\mathcal{A}\pm\mathcal{B}}{2(\varepsilon_{0}+p_{0})(\tau_{q}-\lambda^{\prime})\tau_{\pi}\tau_{\Pi}}\leq 1,}$ (139)
${0\leq\frac{\tau_{q}\left(\gamma^{\prime}\tau_{\pi}+\eta\tau_{\phi}\right)}{(\varepsilon_{0}+p_{0})(\tau_{q}-\lambda^{\prime})\tau_{\pi}\tau_{\phi}}\leq 1,}$ (140)
${0\leq-\frac{\tau_{\tau_{a}}\tilde{\chi}_{2}+\tau_{\tau_{s}}\tilde{\chi}_{3}}{\tau_{\tau_{a}}\tau_{\tau_{s}}}\leq 1,}$ (141)
${0\leq\frac{\tilde{\chi}_{4}}{\tau_{\Theta}}\leq 1.}$ (142)
Under the conditions (III) discussed in Section (III), the coefficients $\tilde{\chi}_{1},\,\tilde{\chi}_{2},\,\tilde{\chi}_{3}$ present in the above conditions are all negative, and $\tilde{\chi}_{4}$ is positive. Additionally, $\tau_{q}>\lambda^{{}^{\prime}}$ as mandated by the stability conditions (125) (132). Since the conditions (138)-(142) all include relaxation times, their values are critical for ensuring causality. Therefore, to illustrate, we will first analyze condition (138). The results obtained will serve as a foundation for analyzing the other conditions. Starting from conditions (III), the coefficients $\tilde{\chi}_{1}$ and $\,\tilde{\chi}_{2}$ are defined as,
${\tilde{\chi}_{1}=\frac{2\beta_{0}}{\chi_{b}}\chi_{1},~{}~{}{\tilde{\chi}_{2}=\frac{\beta_{0}}{\chi_{b}}\chi_{2}},}$
where $\chi_{1},\,\chi_{2}$ are two spin transport coefficients defined in Ref. [28]. The expressions of electric $\chi_{b}$ and magnetic $\chi_{s}$ susceptibilities were explicitly derived in Ref. [6]. Henceforth condition (138) explicitly takes the form,
[…]
where $K_{i}(x)$ is the modified Bessel function of the second type and $x=m/T$ is the ratio of mass to temperature. Clearly, the above condition is satisfied if the relaxation times are sufficiently large. The fulfillment of these conditions allows us to say that second-order dissipative spin hydrodynamics [28] exhibits asymptotic causality; this holds provided the new spin equation of state (3) and sufficiently large relaxation times.
VI Conclusion
@@
In this work, we have analyzed in detail the rest frame stability and asymptotic causality of the second-order spin hydrodynamics. We have used the truncation scheme of the Israel-Stewart formalism derived in our earlier work that extends the minimal causal formulation studied earlier. The crucial property to achieve stability is the use of the spin equation of state, which allows for opposite signs of the electric and magnetic components of the spin density tensor. We demonstrate the rest frame stability of both the first- (Navier-Stokes) and second-order dissipative (Israel-Stewart) spin hydrodynamics in the low-momentum regime. Additionally, the second-order theory exhibits stability and causality at high momentum (with suitably adjusted relaxation times, which is similar to the Israel-Stewart theory).
In our future investigations we plan to study stability and causality in a moving (boosted) frame, as well as the correlation between these two properties. Among the novel approaches that can be used, we are going to consider the information current introduced recently in Refs. [80, 81]. The recent application of this method in minimal spin hydrodynamics [82] makes this approach particularly attractive.
@@
Acknowledgements: A.D. thanks Eduardo Grossi for discussions related to the early stages of this study during the visit to INFN-Florence. A.D. gratefully acknowledges technical discussions with Tomoki Goda, Arpan Das, Shi Pu, David Wagner, and Rajesh Biswas. F.T. acknowledges the hospitality and support of the Institute of Theoretical Physics of the Jagiellonian University. This work was supported in part by the Polish National Agency for Academic Exchange NAWA under the Programme STER–Internationalisation of Doctoral schools, Project no. PPI/STE/2020/1/00020 (A.D.), and the Polish National Science Centre Grants No. 2018/30/E/ST2/00432 (R.R., A.D.) and 2022/47/B/ST2/01372 (W.F.).